A guide of the TMJ style package for L^{A}TEX 2ε

Department of Mathematics Faculty of Science University of Toyama

Jun. 21, 2006

1 Introduction

This TMJ (Toayam Mathematical Journal) (continuation of MJTU(Mathematics Journal of Toyama University)) style file package for L^{A}TEX 2ε will adjust your manuscript into the required form. The usage is very simple. It needs only several commands which are not defined in the standard L^{A}TEX 2ε package. These extra commands will be listed up in the sequel.

Hereafter, we may use the word MJTU instead of TMJ.

2 Installation

First get the TMJ style file package for L^{A}TEX 2ε (which contains this file “guide.tex” ) via WWW from

http://www.sci.u-toyama.ac.jp/math/tmj/index.html

If you get an archive file, you should extract it and put the class file “mjtu2e.cls” into an appropreate directory or folder where your TEX system can search the class file. We show some procedure below depending on which Operating Systems you use.

2.1 Windows/DOS

If you obtained the file “mjtu2e.zip”, extract the zip file using the archiver “pkunzip.exe”. You may obtain this tool via WWW from CTAN (http://www.ctan.org/). After you get these files, put them into an appropreate directory, and in the command prompt type as follows

pkunzip -d mjtu2e.zip

Thereafter put the file mjtu2e.cls into an appropreate directory where your TEX system can search.

2.2 Macintosh

If you obtained the file “mjtu2e.sit.hqx”, extract the file using the archiver “StuffIt Expander”. You can obtain this tool via WWW from Info-Mac

(http://www.info-mac.org/) or from http://www.aladdinsys.com/. Drag

and Drop the file mjtu2e.sit.hqx into the icon StuffIt Expander. And thereafter put the file mjtu2e.cls into an appropreate folder where your TEX system can search.

2.3 UNIX

If you obtained the file “mjtu2e.tar.Z”, extract the file in an appropreate directory, using the programs “uncompress” and “tar”.

cd  ~/temp
cat ~/temp/mjtu2e.tar.Z | uncompress | tar xvf -
cp  mjtu2e.cls  ~/mytexdir/

Usually your TEX system will search the file mjtu2e.cls if you put it into the same directory (~/mytexdir) as the manuscript. If it doesn’t work well, please put the file mjtu2e.cls into the same directory as article.cls, and then update the file “ls-R” (cf. TEX system guide).

3 How to prepare the manuscript

First prepare your manuscript in the standard way for usual L^{A}TEX 2ε files. (Please refer to many good manuals about this).

Thereafter you will make a slight change to reform your manuscript into the required form for MJTU.

3.1 Preambles

Instead of the article.cls file, pleae use the class file “mjtu2e.cls”. For this, you change the first line as follows.

\documentclass{mjtu2e}
%
\title{  }
\author{  }
%
\begin{document}
\maketitle
  body text
\end{document}

Any commands of article.sty (or article.cls) are available.

   • \title{   }
   • \author{   }
   • \thanks{   }

3.1.1 title

The title of the manuscript is printed in 12pt size bold faces.

3.1.2 author

The family names of the authors have to be printed in Small Capitals.

3.1.3 thanks

You may omit this.

3.2 Extra commands in the preamble

For the convenience, the following extra commands are defined in the class file “mjtu2e.cls”.

   • \address{   }
   • \subjclass{   }
   • \keywords{   }
   • \dedicatory{   }

If you want to use these commands, please write them in the following order.

\documentclass{mjtu2e}
\title{      }
\author{     }
\address{    }
\subjclass{  }
\keywords{   }
\dedicatory{   }
\begin{document}
\maketitle
   body text
\end{document}

3.2.1 address

This command is specially defined in mjtu2e.cls. Addresses of authors will be printed in the end of the manuscript. If you have an e-mail address, please write it at the end of each address.

If there are some authors, please separate each address by the command

\sepauthor as follows.
\address{%
Taro Tomiyama\\
Department of Mathematics\\
Faculty of Science\\
University of Toyama\\

Gofuku, Toyama 930-8555, JAPAN\\
e-mail: tomiyama@math.sci.u-toyama.ac.jp\\
\sepauthor
%
Mai de Tsurugi\\
Department of Mathematics\\
Faculty of Science\\
University of Toyama\\
Gofuku, Toyama 930-8555, JAPAN\\
e-mail: tsurugi@math.sci.u-toyama.ac.jp
}

3.2.2 subjclass

This command is specially defined in mjtu2e.cls. You may omit this. By this command, you may indicate “2000 Mathematics Subject Classification”.

3.2.3 keywords

This command is specially defined in mjtu2e.cls. You may omit this. By this command, you may indicate “Key words and phrases”.

3.2.4 dedicatory

As in Ams-L^{A}TEX, you are able to use ”dedicatory” command, which is specially defined in mjtu2e.cls. You may omit this.

3.3 body text

Almost all usual commands may be used. You will prepare the manuscript in the usual way.

3.3.1 abstract

You can write abstract. The abstract is printed in 10pt size. (You may omit this).

\begin{abstract}
\end{abstract}

3.3.2 sections

We may use the following five commands as usual.

   • \section{   }
   • \subsection{   }
   • \subsubsection{   }
   • \paragraph{   }
   • \subparagraph{   }

   After the commands \section{   }, \subsection{   } and \subsubsection{   },

the text will be automatically indented. The period also will be added automatically after each numbering of (sub-)sections.

3.3.3 theorems, propositions, corollaries

In the preamble, define as follows

\newtheorem{thm}{Theorem}
\newtheorem{prop}{Proposition}[section]
\newtheorem{cor}{Corollary}[section]

(Of course you may define them without attaching [section]. You may use appropreate numbering for Theorems, Propositions, etc..)

Then, one may use as follows.

\begin{thm}
\end{thm}

3.3.4 references

Prepare the references as follows.

\begin{thebibliography}{99}
\bibitem{haratomi}K.~Haratomi,
On a topological problem,
{\itshape Japan. J. Math.},
{\bfseries 9}(1932),103-110.
%
\bibitem{kodaira}K.~Kodaira,
{\itshape Complex Manifolds and Deformation of Complex Structures},
Grundlehren der mathematischen Wissenschaften 283,
Springer-Verlag, New York, Berlin, Heidelberg, Tokyo, 1986.
%
\bibitem{milnor}J.~Milnor,
{\itshape Morse Theory},
Annals of Mathematics Studies 51,
Princeton Univ. Press, Princeton, NJ, 1963.
%
\end{thebibliography}

4 Check the manuscript

Finally please check that your manuscript satisfies the following required form.

1. textheight = 195 mm, textwidth = 125 mm.

2. A space 50 mm deep should be left at the top of the first page.

3. Title of Paper in 12pt size boldface, centered and not underlined.

4.

Author’s name: First name(s) of which the first letter in capital and remains in lowercases, followed by Surname of which the first letter in capital and remains in small capitals, without the preposition “by”.

5. Abstract (may be omitted).

6. Table of Contents (required only for very long papers).

7. Text (in 11pt size).

8. References.

9.

Author’s full postal address with zip code aligned on the left under the last reference in the bibliography. Where possible, e-mail addresses should be given as well.

Department of Mathematics Faculty of Science University of Toyama Gofuku, Toyama 930-8555, JAPAN